OpenStudy (anonymous):

The number of days needed to remodel a house varies inversely with the number of people working on the job. It takes 18 weeks for 4 people to complete the project. If the job has to be finished in 8 weeks, how many people are needed?

OpenStudy (anonymous):

Let 'd' be the number of days needed and 'p' be the number of people needed. d∝1/p d=k/p If it takes 18 weeks (126 days) for 4 people to complete project... 126=k/4 k=126*4 =504 Therefore your equation is d=504/p If the job has to be finished in 8 weeks (56 days), how many people are needed? Subst d=56 into d=504/p and solve for p 56=504/p 56p=504 p=504/56 = 9 9 people are needed. Do you follow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i have another one im gonna post right here

OpenStudy (anonymous):

A rectangle has length 36 cm and width 28 cm. Find the length of another rectangle of equal area whose width is 21 cm.

OpenStudy (anonymous):

A=length x width = 36 * 28 =1008 cm^2 We need the length of a rectangle with a width of 21 cm and an area of 1008 cm^2 A= length x width 1008 = length * 21 1008/21 = length 48 = length The length would be 48 cm

OpenStudy (anonymous):

an easier way-the area is =36*28 =21*length length=36*28/21=48cm=length